# Neuromorphic quadratic programming for efficient and scalable model predictive control

### v1

@@ Abstract
Convex QP problems arising in size-, weight-, and power-constrained (SWaP) systems are conventionally solved on CPUs, even though a few solutions have been developed to leverage the parallel compute capabilities of GPUs, FPGAs, or ASICs. A range of high-performance QP solvers exist for CPUs, such as GUROBI [10], MOSEK [11] and CVXOPT [12]. Lightweight CPU solvers specifically optimized for embedded systems include qpOASES [13], ECOS [14] and OSQP [15]; such solvers avoid any dependence on large external libraries, use only basic operations (e.g. avoiding division), minimize the steps to solution, optimize the code for mobile processors, and often parallelize their code. Unfortunately, CPUs in general — and the ones in embedded systems in particular — often do not support the degree of parallelism needed to accelerate large optimization workloads.
@@ II-B Neuromorphic hardware implementation
For QPs arising in MPC, the neuromorphic architecture of Loihi 2 already offers advantages over conventional architectures and ANN accelerators because of the compute and memory co-location which will be further discussed in section III. However, first-order type algorithms like those in eq. (4)-(8) can be made more suitable for neuromorphic hardware by employing - coding which can be implemented in Loihi 2 [24] . Doing so sparsifies the spiking activity at the cost of solution accuracy. This,in turn, improves the time and energy to solution because of fewer messages/spikes being moved around and, as a consequence, fewer operations in general. However, this reduces operations within acceptable solution degradation only when the datatypes on chip are more precise (16-/32-bit floating-point) which is currently not the case in Loihi 2. Therefore, in this article we focus on results obtained without employing sigma-delta coding.
@@ III Quadratic programming for model predictive control
III-A Problem definition
To demonstrate the value of our neuromorphic QP solver, we apply it to solving the QPs arising in model predictive control in the quadrupedal robot ANYmal [9], a state-of-the-art quadrupedal robot. At the core of an MPC iteration is the mathematical optimization problem of minimizing the error between actual and goal trajectories, i.e., the tracking error. The optimization is subject to constraints, such as the limits posed by the rigid-body dynamics, joints, actuators, and environment. As such, the optimization problem is a computationally expensive, nonlinear (and perhaps non-convex). However, it is typically ‘linearized-and-quadratized’ using techniques akin to Taylor series expansion, resulting in a convex QP problem from eq. (1) and (2). MPC is a computationally intensive control scheme that must be executed under strict time budgets in real-time control loops. Solving the QP accounts for a major chunk of computational time in MPC of ANYmal. The QP further increases in complexity with larger time horizons covered by the MPC or more degrees of freedom of the robot. In order to meet these requirements, MPC is solved on a dedicated laptop-class CPU. Running the QP in the MPC on Loihi would be the first step towards making computation for control more suitable for SWaP systems.
@@ III-A Problem definition
III-B Neural and synaptic scaling
@@
III-C Benchmarking procedure and dataset
The efficiency, scalability, and speed of the Loihi QP solver was benchmarked against OSQP running on a standard CPU and an embedded CPU, and against the CUDA-optimized version of it, cuOSQP, running on a GPU 333Loihi 2: Oheo Gulch board running Lava v0.7.0 and Lava-Optimization v0.2.4 with an Intel Core i7-9700K CPU host with 32GB RAM running Ubuntu 20.04.5 LTS.
@@ III-C Benchmarking procedure and dataset
Evaluation metrics were time to solution (TTS), energy to solution (ETS) and energy-delay product (EDP) for different problem sizes. OSQP is a CPU-based state-of-the-art solver for convex QP problems with linear inequality constraints [15].
The dataset used in this paper consists of data from 2173 individual time-steps of MPC of an ANYmal robot [25]. The dataset consists of the matrices obtained after linearization and quadratization of a nonlinear objective function and penalizes deviation from a reference trajectory as well as the ANYmal dynamics. We tile these matrices in an appropriate manner to construct the matrix of the QP objective function from (1) and the matrix of the constraints from (2). This tiling is demonstrated in fig. 4. The derivation and explanation of this tiling has been omitted here since it’s beyond the scope of the article. It can be seen that these matrices are mostly block-diagonally populated and very sparse. This entails sparse matrix-vector multiplication making a suitable candidate for Loihi 2. The vectors and from the cost and constraints respectively are constructed by stacking the vectors associated with the tracking cost and the dynamics in each stage of the MPC. We investigate how well the performance of the solver scales with problem size by varying the number of variables. In general, the problem size is determined by the time horizon and the degrees of freedom of the robot. Here, we choose six horizon lengths time-steps, resulting in problems with variables, respectively. We use this as a proxy for a robot with higher degrees of freedom. Most QPs in the dataset are similar in difficulty (similar condition numbers). Therefore, for each horizon length, we have chosen 10 different QPs of varying difficulty that best represent all the problems in the dataset. In summary, we have a dataset containing a total of 60 representative QP problems of 6 different sizes from the ANYmal dataset, spanning from to variables.
III-D Convergence, power and performance analysis
@@
Further Fig. 6c shows a energy-delay product advantage for the largest problem size when using Loihi 2 in place of a laptop CPU. It can be seen in fig. 6c that the EDP curves for CPUs (embedded and laptop) closely resemble each other. The laptop and embedded CPU thus trade TTS against ETS, but the combined EDP is characteristic for CPUs. For a GPU, both the ETS and EDP are orders of magnitude higher mainly.
@@ IV Conclusion
We have developed an efficient and scalable approach for solving convex quadratic constraint optimization problems with neuromorphic chips. The benefits of neuromorphic systems can be exploited for optimization algorithms by interpreting the structure of many iterative optimization approaches—such as gradient descent and primal-dual updates—as the dynamics of recurrent neural networks. To demonstrate the benefits of this framework, we implemented a QP solver on Intel’s Loihi 2 research chip and made it publicly available as part of the Lava software framework. We apply it to solving QPs on a dataset extracted from the model predictive control (MPC) pipeline of the ANYmal quadrupedal robot. Our Loihi solver achieves similar speed as the state-of-the-art solver OSQP on a modern CPU [28], which was the best von-Neumann-based-candidate for comparison, for approximate solutions while consuming about two orders of magnitude less energy. Most notably, time and energy to solution on Loihi scale better with increasing problem complexity than OSQP on CPU.
@@
The most significant limitation of the current approach is the limited bit-precision of state variables and synaptic connections on neuromorphic architectures like Loihi. This can lead to lower solution optimality than solvers on conventional floating-point architectures. Such limitations can be circumvented by allocating multiple low-bit variables to effectively achieve higher precision arithmetic or enabling floating-point support in future chip generations. Nonetheless, preliminary observations in closed-loop control simulations hint that fast approximate solutions are sufficient for iterative scenarios like MPC with sequential warm starts. We invite robotics labs interested in assessing the performance of our fixed-point Quadratic Programming (QP) solver on their cyber-physical systems to become part of Intel’s Neuromorphic Research Community (INRC). By joining INRC, labs will gain access to the proprietary code for running the solver on Loihi 2. They can then apply for borrowing a Loihi 2 board for closed-loop testing in their MPC applications. Further, with floating-point support, the current gains observed with compute-memory co-location can be bolstered by employing sigma-delta coding which would reduce spiking activity in the chip and consequently TTS and ETS.

### v2

@@ Abstract
Convex QP problems arising in size-, weight-, and power-constrained (SWaP) systems are conventionally solved on CPUs, even though a few solutions have been developed to leverage the parallel compute capabilities of GPUs, FPGAs, or ASICs. A range of high-performance QP solvers exist for CPUs, such as GUROBI [10], MOSEK [11], SCS [12], and CVXOPT [13]. Lightweight CPU solvers specifically optimized for embedded systems include qpOASES [14], ECOS [15] and OSQP [16]; such solvers avoid any dependence on large external libraries, use only basic operations (e.g. avoiding division), minimize the steps to solution, optimize the code for mobile processors, or often parallelize their code. Unfortunately, CPUs in general — and the ones in embedded systems in particular — often do not support the degree of parallelism needed to accelerate large optimization workloads.
@@ II-B Neuromorphic hardware implementation
For QPs arising in MPC, the neuromorphic architecture of Loihi 2 already offers advantages over conventional architectures and ANN accelerators because of the compute and memory co-location which will be further discussed in section III. However, first-order type algorithms like those in eq. (4)-(8) can be made more suitable for neuromorphic hardware by employing - coding which can be implemented in Loihi 2 [25] . Doing so sparsifies the spiking activity at the cost of solution accuracy. This, in turn, improves the time and energy to solution because of fewer messages/spikes being moved around and, as a consequence, fewer operations in general. However, this reduces operations within acceptable solution degradation only when the datatypes on chip are more precise (16-/32-bit datatypes) which is currently not the case in Loihi 2. Therefore, in this article we focus on results obtained without employing sigma-delta coding.
@@ III Quadratic programming for model predictive control
To demonstrate the value of our neuromorphic QP solver, we apply it to solving the QPs arising in model predictive control of a state-of-the-art quadrupedal ANYmal robot [9].
III-A The ANYmal robot
The ANYmal platform comprises a series of quadrupedal robots designed for different kinds of tasks like inspection, surveillance, and search and rescue missions. The robots are equipped with a suite of sensors enabling autonomous navigation and perception of their surroundings, as well as a comprehensive software ecosystem. The robot is commercially available and deployed in a range of industrial and commercial settings. However, model predictive control for this robot has on the order of variables. Due to this high dimensionality, each robot requires a separate Intel® Core™ i7 processor in order to solve model predictive control within acceptable time budgets. Further, as a mobile platform, bringing down power utilization would contribute to longer uptimes. Loihi could be used as a co-processor to solve model predictive control with very little additional burden on the battery. We therefore use data acquired from a physical ANYmal robot performing tasks provided by the ANYmal team.
III-B Problem definition
At the core of an MPC iteration is the mathematical optimization problem of minimizing the error between actual and goal trajectories, i.e., the tracking error. The optimization is subject to constraints, such as the limits posed by the rigid-body dynamics, joints, actuators, and environment. As such, the optimization problem is a computationally expensive, nonlinear (and perhaps non-convex). However, it is typically ‘linearized-and-quadratized’ using techniques akin to Taylor series expansion, resulting in a convex QP problem from eq. (1) and (2). MPC is a computationally intensive control scheme that must be executed under strict time budgets in real-time control loops. Solving the QP accounts for a major chunk of computational time in MPC of ANYmal. The QP further increases in complexity with larger time horizons covered by the MPC or more degrees of freedom of the robot. In order to meet these requirements, MPC is solved on a dedicated laptop-class CPU. Running the QP in the MPC on Loihi would be the first step towards making computation for control more suitable for SWaP systems.
@@ III-B Problem definition
III-C Neural and synaptic scaling
@@
III-D Benchmarking procedure and dataset
The efficiency, scalability, and speed of the Loihi QP solver was benchmarked against 1) CVXOPT, 2) SCS solver running on a laptop CPU and OSQP running on 3) a standard CPU, 4) an embedded CPU, and 5) against the CUDA-optimized version of it, cuOSQP, running on a GPU. 333Loihi 2: Oheo Gulch board running Lava v0.7.0 and Lava-Optimization v0.2.4 with an Intel Core i7-9700K CPU host with 32GB RAM running Ubuntu 20.04.5 LTS.
@@ III-D Benchmarking procedure and dataset
Evaluation metrics were time to solution (TTS), energy to solution (ETS) and energy-delay product (EDP) for different problem sizes. OSQP is a CPU-based state-of-the-art solver for convex QP problems with linear inequality constraints [16]. We ultimately use the OSQP solver as a reference for our Loihi-based solver since it was the best performing CPU solver.
The dataset used in this paper consists of data from 2173 individual time-steps of MPC of an ANYmal robot [26]. The dataset consists of the matrices obtained after linearization and quadratization of a non-linear objective function and penalizes deviation from a reference trajectory as well as the ANYmal dynamics. We tile these matrices in an appropriate manner to construct the matrix of the QP objective function from (1) and the matrix of the constraints from (2). This tiling is demonstrated in fig. 4. The derivation and explanation of this tiling has been omitted here since it’s beyond the scope of the article. It can be seen that these matrices are mostly block-diagonally populated and very sparse. This entails sparse matrix-vector multiplication making a suitable candidate for Loihi 2. The vectors and from the cost and constraints respectively are constructed by stacking the vectors associated with the tracking cost and the dynamics in each stage of the MPC. We investigate how well the performance of the solver scales with problem size by varying the number of variables. In general, the problem size is determined by the time horizon and the degrees of freedom of the robot. Here, we choose six horizon lengths time-steps, resulting in problems with variables, respectively. We use this as a proxy for a robot with higher degrees of freedom. Most QPs in the dataset are similar in difficulty (similar condition numbers). Therefore, for each horizon length, we have chosen 10 different QPs of varying difficulty that best represent all the problems in the dataset. In summary, we have a dataset containing a total of 60 representative QP problems of 6 different sizes from the ANYmal dataset, spanning from to variables.
III-E Convergence, power and performance analysis
@@
Further, fig. 6c shows a energy-delay product advantage for the largest problem size when using Loihi 2 in place of a laptop CPU. It can be seen in fig. 6c that the EDP curves for CPUs (embedded and laptop) closely resemble each other. The laptop and embedded CPU thus trade TTS against ETS, but the combined EDP is characteristic for CPUs. For a GPU, both the ETS and EDP are orders of magnitude higher mainly.
@@ IV Conclusion
We have developed an efficient and scalable approach for solving convex quadratic constraint optimization problems with neuromorphic chips. The benefits of neuromorphic systems can be exploited for optimization algorithms by interpreting the structure of many iterative optimization approaches—such as gradient descent and primal-dual updates—as the dynamics of recurrent neural networks. To demonstrate the benefits of this framework, we implemented a QP solver on Intel’s Loihi 2 research chip and made it publicly available as part of the Lava software framework. To illustrate its benefits, we applied the QP solver to tackle workloads that arose in the model predictive control pipeline of ANYmal during its normal operation. Our Loihi solver solved the workloads with similar speed as the state-of-the-art solver OSQP on a modern CPU [29], which was the best von-Neumann-based-candidate for comparison, for approximate solutions while consuming about two orders of magnitude less energy. Most notably, time and energy to solution on Loihi scale better with increasing problem complexity than OSQP on CPU.
@@
The most significant limitation of the current approach is the limited bit-precision of state variables and synaptic connections on neuromorphic architectures like Loihi. This can lead to lower solution optimality than solvers on conventional floating-point architectures. Nevertheless, our solver consistently achieved solutions that deviated by less than from the true optimal solution for QPs extracted during a real-world operation of a physical ANYmal robot. For applications that require a higher precision, the limited precision of the Loihi solver can be circumvented by allocating multiple low-bit variables to effectively achieve higher precision arithmetic or enabling higher precision arithmetic in general in future chip generations. Nonetheless, preliminary observations in closed-loop control simulations of ANYmal hint that fast approximate solutions are sufficient for iterative scenarios like MPC with sequential warm starts. We hypothesize this is partly because the result of each MPC cycle is only used to determine motor commands for the next few time steps, despite optimizing over time horizons of more than 100 time steps. After controlling the next few steps, new sensory recordings are taken and the next MPC cycle can iteratively correct any errors resulting from, e.g., the linearized mechanical robotic model, environmental perturbations, and limited bit-precision. We invite robotics labs interested in assessing the performance of our fixed-point QP solver on their cyber-physical systems to become part of Intel’s Neuromorphic Research Community (INRC). By joining the INRC, labs will gain access to the proprietary code for running the solver on Loihi 2. They can then apply for borrowing a Loihi 2 board for closed-loop testing in their MPC applications. The implementation of our solver in Lava, a software framework built to support asynchronous message passing, bodes well for integration with the asynchronous publish and subscribe system of ROS. Further, with floating-point support, the current gains observed with compute-memory co-location can be bolstered by employing sigma-delta coding which would reduce spiking activity in the chip and consequently TTS and ETS.